Town A and B are 390 km apart. A car travels from town A at 70 km/h. At the same time
a Van starts to travel at 60km/hr from Town B. When and where do the car and the van
meet each other?
1
Expert's answer
2021-07-26T00:53:01-0400

"Speed =\frac{Distance}{Time}"

"Time=\frac{Distance}{Time}"

Distance=390Km

Speed of car=70km/h

Speed of van =60km/h

Since the two vehicles are travelling at different directions;

The total speed = 70+60=130km/h

"Time=\frac{390km}{130km/h}=3hrs"

The and the van will meet after 3hrs

The two vehicles will met at

"Distance=speed\u00d7time"

Meeting point =70km/h×3h=210km

The meeting point will be 210km from town A